# Scale-Invariant Gradient Aggregation for Constrained Multi-Objective Reinforcement Learning

### v1

@@ A.2 Proof of Theorems
We first check the constraint. Due to the $L$-Lipschitz continuity assumption,
$J_{C_{k}}(\theta_{\mathrm{old}}+\beta\bar{g}_{\omega}^{\mathrm{ag}})-J_{C_{k}}(\theta_{\mathrm{old}})\leq\beta b_{k}^{T}\bar{g}_{\omega}^{\mathrm{ag}}+\frac{L}{2}\beta^{2}||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2}.$ (21)
From Equation (8), $b_{k}^{T}\bar{g}_{\omega}^{\mathrm{ag}}\leq d_{k}-K\epsilon-J_{C_{k}}(\theta_{\mathrm{old}})$ for $k\in I_{\mathrm{AC}}^{\epsilon}$. Reordering the above inequality,
$J_{C_{k}}(\theta_{\mathrm{old}}+\beta\bar{g}_{\omega}^{\mathrm{ag}})-d_{k}\leq(1-\beta)(J_{C_{k}}(\theta_{\mathrm{old}})-d_{k})+\beta(\frac{L}{2}\beta||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2}-K\epsilon)\leq 0,$ (22)
since $0\leq\beta\leq 1$ and $\beta\leq 2K\epsilon/(L||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2})$. Thus, the constraint is also satisfied at $\theta_{\mathrm{new}}$. Subsequently, we check the objective for $i\in I_{\omega}$. Using the $L$-Lipschitz continuity property and (8),
$J_{R_{i}}(\theta_{\mathrm{old}}+\beta\bar{g}_{\omega}^{\mathrm{ag}})-J_{R_{i}}(\theta_{\mathrm{old}})\geq\beta g_{i}^{T}\bar{g}_{\omega}^{\mathrm{ag}}-\frac{L}{2}\beta^{2}||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2}\geq\beta(e_{i}-\frac{L}{2}\beta||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2}).$ (23)
Since $\beta\leq 2||\bar{g}_{\omega}^{\mathrm{ag}}||_{H}/(\alpha L||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2})\leq 2e_{i}/(L||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2})$,
$J_{R_{i}}(\theta_{\mathrm{old}}+\beta\bar{g}_{\omega}^{\mathrm{ag}})-J_{R_{i}}(\theta_{\mathrm{old}})\geq 0.$ (24)
As a result, updating the policy parameters to $\theta_{\mathrm{new}}$ will increase the objective for $i\in I_{\omega}$. ∎
@@
Since $g_{\omega}^{\mathrm{ag}}=\beta_{1}\beta_{2}\bar{g}_{\omega}^{\mathrm{ag}}$, where $\beta_{1}=\mathrm{min}(1,\epsilon/||\bar{g}_{\omega}^{\mathrm{ag}}||_{H})$ and $\beta_{2}:=\mathrm{min}(1,G/\alpha)$, we only need to show that the step size $\beta=\beta_{1}\beta_{2}$ satisfies the condition of Lemma 4.1 and that $J_{C_{k}}(\theta_{\mathrm{new}})\leq d_{k}$ for $\forall k$. There are two cases regarding $\alpha$: 1) $\alpha\leq G$ and 2) $\alpha>G$. For the fist case $\alpha\leq G$, $\beta=\beta_{1}$ since $\beta_{2}=1$. Since $0\leq\beta_{1}\leq 1$, it is obvious that $0\leq\beta\leq 1$. Thus, we only need to check $1\leq 2||\bar{g}_{\omega}^{\mathrm{ag}}||_{H}/(\alpha\beta L||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2})$. Since the matrix $H$ is positive definite, there exist $h>0$ such that $h||x||_{2}\leq||x||_{H}$. Then, the following inequality holds:
$\frac{2||\bar{g}_{\omega}^{\mathrm{ag}}||_{H}}{\alpha L\beta||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2}}\geq\frac{2h^{2}}{\alpha L\beta_{1}||\bar{g}_{\omega}^{\mathrm{ag}}||_{H}}\geq\frac{2h^{2}}{\alpha L\epsilon}\geq\frac{2h^{2}}{GL\epsilon}.$ (25)
Thus, if $\epsilon\leq 2h^{2}/GL$, the step size satisfies Lemma 4.1. For the second case $\alpha>G$, $\beta=G\beta_{1}/\alpha$. Similar to the first case, the following inequality holds:
$\frac{2||\bar{g}_{\omega}^{\mathrm{ag}}||_{H}}{\alpha L\beta||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2}}\geq\frac{2h^{2}}{GL\beta_{1}||\bar{g}_{\omega}^{\mathrm{ag}}||_{H}}\geq\frac{2h^{2}}{GL\epsilon}.$ (26)
Therefore, if $\epsilon\leq 2h^{2}/GL$, the step size condition for Lemma 4.1 is satisfied.
Second, we have to check that $J_{C_{k}}(\theta_{\mathrm{new}})\leq d_{k}$ for $k\notin I_{\mathrm{AC}}^{\epsilon}$. Since $k\notin I_{\mathrm{AC}}^{\epsilon}$, $J_{C_{k}}(\theta_{\mathrm{old}})<d_{k}-\epsilon(K+\sqrt{b_{k}^{T}H^{-1}b_{k}})$. Additionally, from the result of Appendix A.1, we can get the maximum of $b_{k}^{T}g$ in the local region by substituting the optimal $g^{*}$ into $b_{k}^{T}g$, which results in $\epsilon\sqrt{b_{k}^{T}H^{-1}b_{k}}$. This means $b_{k}^{T}g_{\omega}^{\mathrm{ag}}\leq\epsilon\sqrt{b_{k}^{T}H^{-1}b_{k}}$, since $g_{\omega}^{\mathrm{ag}}$ is within the local region. Using the above facts and the $L$-Lipschitz continuity assumption,
$J_{C_{k}}(\theta_{\mathrm{old}}+\beta_{1}\beta_{2}\bar{g}_{\omega}^{\mathrm{ag}})$ $\leq\beta_{1}\beta_{2}b_{k}^{T}\bar{g}_{\omega}^{\mathrm{ag}}+\frac{L}{2}\beta_{1}^{2}\beta_{2}^{2}||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2}+J_{C_{k}}(\theta_{\mathrm{old}})$ (27)
$\leq\beta b_{k}^{T}\bar{g}_{\omega}^{\mathrm{ag}}+\frac{L}{2}\epsilon^{2}||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2}/||\bar{g}_{\omega}^{\mathrm{ag}}||_{H}^{2}+J_{C_{k}}(\theta_{\mathrm{old}})$
$\leq\beta b_{k}^{T}\bar{g}_{\omega}^{\mathrm{ag}}+\frac{L}{2h^{2}}\epsilon^{2}+J_{C_{k}}(\theta_{\mathrm{old}})$
$\leq\epsilon\sqrt{b_{k}^{T}H^{-1}b_{k}}+\frac{L}{2h^{2}}\epsilon^{2}+J_{C_{k}}(\theta_{\mathrm{old}})\leq\frac{L}{2h^{2}}\epsilon^{2}-K\epsilon+d_{k}.$
Here, if $\epsilon\leq(2Kh^{2})/L$, $J_{C_{k}}(\theta_{\mathrm{new}})\leq d_{k}$. As a result, if $\epsilon\leq 2h^{2}\min(K,1/G)/L$, CoMOGA improves the objectives for $i\in I_{\omega}$ and satisfies $J_{C_{k}}(\theta_{\mathrm{new}})\leq d_{k}$ for $\forall k$. ∎
See 4.5

### v2

@@ B.1 Proof of Transformation Process
Let us assume that there exists a policy $\mu$ that dominates the LS optimal policy $\pi_{\omega}^{*}$. By definition of the constrained dominance, the following is satisfied:
$J_{R_{i}}(\mu)>J_{R_{i}}(\pi_{\omega}^{*})\;\forall i\;\text{and}\;J_{C_{k}}(\mu)\leq d_{k}\;\forall k.\Rightarrow\sum_{i=1}^{N}\omega_{i}(J_{R_{i}}(\mu)-J_{R_{i}}(\pi_{\omega}^{*}))>0.$
However, this equation contradicts the fact that $\pi_{\omega}^{*}$ is an LS optimal policy. Consequently, $\pi_{\omega}^{*}$ is not dominated by any others, it is also a CP optimal policy. ∎
@@
We first consider the case $J_{C_{k}}(\theta_{t})\leq d_{k}\;\forall k$, in which the policy is updated as follows:
$\theta_{t+1}=\theta_{t}+\frac{\alpha_{t}}{1-\gamma}(\sum_{i=1}^{N}\nu_{t,i}^{a}A_{R_{i}}^{t}-\alpha_{t}\sum_{k=1}^{M}\lambda_{t,k}^{a}A_{C_{k}}^{t}).$
Using Lemma 1 in [29], we can derive the following inequality:
$\sum_{i}\nu_{t,i}^{a}(J_{R_{i}}(\theta_{t+1})-J_{R_{i}}(\theta_{t}))-\alpha_{t}\sum_{k}\lambda_{t,k}^{a}(J_{C_{k}}(\theta_{t+1})-J_{C_{k}}(\theta_{t}))$ (17)
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{t+1}}\left[\mathbb{E}_{a\sim\pi_{t+1}(\cdot|s)}\left[\sum_{i=1}^{N}\nu_{t,i}^{a}A_{R_{i}}^{t}(s,a)-\alpha_{t}\sum_{k=1}^{M}\lambda_{t,k}^{a}A_{C_{k}}^{t}(s,a)\right]\right]$
$=\frac{1}{\alpha_{t}}\mathbb{E}_{s\sim d_{\rho}^{t+1}}\left[\mathbb{E}_{a\sim\pi_{t+1}(\cdot|s)}\left[\theta_{t+1}(s,a)-\theta_{t}(s,a)\right]\right]$
$=\frac{1}{\alpha_{t}}\mathbb{E}_{s\sim d_{\rho}^{t+1}}\left[\mathbb{E}_{a\sim\pi_{t+1}(\cdot|s)}\left[\log\frac{\pi_{t+1}(a|s)Z_{t}(s)}{\pi_{t}(a|s)}\right]\right]$
$=\frac{1}{\alpha_{t}}\mathbb{E}_{s\sim d_{\rho}^{t+1}}\left[D_{\mathrm{KL}}(\pi_{t+1}(\cdot|s)||\pi_{t}(\cdot|s))+\log Z_{t}(s)\right]$
[…]
where (i) follows from the fact that $d_{\rho}^{\pi}=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathrm{Pr}(s_{t}=s)\geq(1-\gamma)\rho(s)$. By Lemma 5 in [33],
$|J_{R_{i}}(\theta_{t+1})-J_{R_{i}}(\theta_{t})|$ $\leq\frac{2R_{\mathrm{max}}}{1-\gamma}||\theta_{t+1}-\theta_{t}||_{2}\;\forall i,$
$|J_{C_{k}}(\theta_{t+1})-J_{C_{k}}(\theta_{t})|$ $\leq\frac{2R_{\mathrm{max}}}{1-\gamma}||\theta_{t+1}-\theta_{t}||_{2}\;\forall k.$
$\Rightarrow$ $\Big{|}\sum_{i}\nu_{t,i}^{a}(J_{R_{i}}(\theta_{t+1})-J_{R_{i}}(\theta_{t}))-\alpha_{t}\sum_{k}\lambda_{t,k}^{a}(J_{C_{k}}(\theta_{t+1})-J_{C_{k}}(\theta_{t}))\Big{|}$
$\leq\frac{2R_{\mathrm{max}}}{1-\gamma}(N+\alpha_{t}M)\lambda_{\max}||\theta_{t+1}-\theta_{t}||_{2}$
$=\frac{2R_{\mathrm{max}}}{1-\gamma}(N+\alpha_{t}M)\lambda_{\max}\Big{|}\Big{|}\frac{\alpha_{t}}{1-\gamma}(\sum_{i=1}^{N}\nu_{t,i}^{a}A_{R_{i}}^{t}-\alpha_{t}\sum_{k=1}^{M}\lambda_{t,k}^{a}A_{C_{k}}^{t})\Big{|}\Big{|}_{2}$
$\leq\frac{2\alpha_{t}R^{2}_{\mathrm{max}}}{(1-\gamma)^{3}}(N+\alpha_{t}M)^{2}\lambda_{\max}^{2}\sqrt{|S||A|}.$
Then, (17) can be rewritten as follows:
$\frac{2\alpha_{t}R^{2}_{\mathrm{max}}}{(1-\gamma)^{3}}(N+\alpha_{t}M)^{2}\lambda_{\max}^{2}\sqrt{|S||A|}\geq\frac{1-\gamma}{\alpha_{t}}\mathbb{E}_{s\sim\rho}\left[\log Z_{t}(s)\right],$ (18)
which is satisfied for any $\rho$. Finally, we derive the following inequality for a policy $\mu$ as follows:
$\sum_{i}\nu_{t,i}^{a}(J_{R_{i}}(\mu)-J_{R_{i}}(\theta_{t}))-\alpha_{t}\sum_{k}\lambda_{t,k}^{a}(J_{C_{k}}(\mu)-J_{C_{k}}(\theta_{t}))$ (19)
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\mu}}\left[\mathbb{E}_{a\sim\mu(\cdot|s)}\left[\sum_{i=1}^{N}\nu_{t,i}^{a}A_{R_{i}}^{t}(s,a)-\alpha_{t}\sum_{k=1}^{M}\lambda_{t,k}^{a}A_{C_{k}}^{t}(s,a)\right]\right]$
$=\frac{1}{\alpha_{t}}\mathbb{E}_{s\sim d_{\rho}^{\mu}}\left[\mathbb{E}_{a\sim\mu(\cdot|s)}\left[\log\frac{\pi_{t+1}(a|s)Z_{t}(s)}{\pi_{t}(a|s)}\right]\right]$
$=\frac{1}{\alpha_{t}}\mathbb{E}_{s\sim d_{\rho}^{\mu}}\left[D_{\mathrm{KL}}(\mu(\cdot|s)||\pi_{t}(\cdot|s))-D_{\mathrm{KL}}(\mu(\cdot|s)||\pi_{t+1}(\cdot|s))+\log Z_{t}(s)\right]$
[…]
$\quad+\frac{2\alpha_{t}R^{2}_{\mathrm{max}}}{(1-\gamma)^{3}}(N+\alpha_{t}M)^{2}\lambda_{\max}^{2}\sqrt{|S||A|},$
where (i) is from (18) by replacing $\rho$ with $d_{\rho}^{\mu}$. For brevity, we denote $\mathbb{E}_{s\sim d_{\rho}^{\mu}}\left[D_{\mathrm{KL}}(\mu(\cdot|s)||\pi_{t}(\cdot|s))\right]$ as $D_{\mathrm{KL}}(\mu||\pi_{t})$. Now, we consider the second case where the constraints are violated. In this case,
$\theta_{t+1}=\theta_{t}+\frac{\alpha_{t}}{1-\gamma}(\alpha_{t}\sum_{i=1}^{N}\nu_{t,i}^{b}A_{R_{i}}^{t}-\sum_{k=1}^{M}\lambda_{t,k}^{b}A_{C_{k}}^{t}),$
which is symmetrical to the first case, achieved by replacing $\nu_{t,i}^{a}\rightarrow\alpha_{t}\nu_{t,i}^{b}$ and $\alpha_{t}\lambda_{t,k}^{a}\rightarrow\lambda_{t,k}^{b}$. Using this symmetry property, we can obtain the following inequality from (19):
$\alpha_{t}\sum_{i}\nu_{t,i}^{b}(J_{R_{i}}(\mu)-J_{R_{i}}(\theta_{t}))-\sum_{k}\lambda_{t,k}^{b}(J_{C_{k}}(\mu)-J_{C_{k}}(\theta_{t}))$ (20)
$\leq\frac{1}{\alpha_{t}}(D_{\mathrm{KL}}(\mu||\pi_{t})-D_{\mathrm{KL}}(\mu||\pi_{t+1}))+\frac{2\alpha_{t}R^{2}_{\mathrm{max}}}{(1-\gamma)^{3}}(\alpha_{t}N+M)^{2}\lambda_{\max}^{2}\sqrt{|S||A|}.$
Now, we define a set of time steps, $\mathcal{N}:=\{t|J_{C_{k}}(\pi_{t})\leq d_{k}\;\forall k\}$. Then, by summing (19) and (20) over several time steps, the following is satisfied for a policy $\mu$:
$\sum_{t\in\mathcal{N}}\left(\alpha_{t}\sum_{i}\nu_{t,i}^{a}(J_{R_{i}}(\mu)-J_{R_{i}}(\theta_{t}))-\alpha_{t}^{2}\sum_{k}\lambda_{t,k}^{a}(J_{C_{k}}(\mu)-J_{C_{k}}(\theta_{t}))\right)$ (21)
$+\sum_{t\notin\mathcal{N}}\left(\alpha_{t}^{2}\sum_{i}\nu_{t,i}^{b}(J_{R_{i}}(\mu)-J_{R_{i}}(\theta_{t}))-\alpha_{t}\sum_{k}\lambda_{t,k}^{b}(J_{C_{k}}(\mu)-J_{C_{k}}(\theta_{t}))\right)$
$\leq D_{\mathrm{KL}}(\mu||\pi_{0})+\sum_{t}\alpha_{t}^{2}\underbrace{\frac{2R^{2}_{\mathrm{max}}}{(1-\gamma)^{3}}\max(\alpha_{0},1)^{2}(N+M)^{2}\lambda_{\max}^{2}\sqrt{|S||A|}}_{=:\kappa_{1}}.$
If $t\notin\mathcal{N}$, $\lambda_{t,k}^{b}=0$ for $J_{C_{k}}(\theta_{t})\leq d_{k}$, which results in $\sum_{k}\lambda_{t,k}^{b}(J_{C_{k}}(\pi_{f})-J_{C_{k}}(\theta_{t}))\leq-\eta$. By using this fact and replacing $\mu$ in (21) with $\pi_{f}$,
$\sum_{t\in\mathcal{N}}\left(\alpha_{t}\sum_{i}\nu_{t,i}^{a}(J_{R_{i}}(\pi_{f})-J_{R_{i}}(\theta_{t}))-\alpha_{t}^{2}\sum_{k}\lambda_{t,k}^{a}(J_{C_{k}}(\pi_{f})-J_{C_{k}}(\theta_{t}))\right)$
$+\sum_{t\notin\mathcal{N}}\left(\alpha_{t}^{2}\sum_{i}\nu_{t,i}^{b}(J_{R_{i}}(\pi_{f})-J_{R_{i}}(\theta_{t}))+\alpha_{t}\eta\right)\leq D_{\mathrm{KL}}(\pi_{f}||\pi_{0})+\kappa_{1}\sum_{t}\alpha_{t}^{2}.$
$\Rightarrow\sum_{t\in\mathcal{N}}\left(\alpha_{t}\sum_{i}\nu_{t,i}^{a}(J_{R_{i}}(\pi_{f})-J_{R_{i}}(\theta_{t}))\right)+\sum_{t\notin\mathcal{N}}\alpha_{t}\eta$
$\leq D_{\mathrm{KL}}(\pi_{f}||\pi_{0})+\kappa_{1}\sum_{t}\alpha_{t}^{2}+\sum_{t\in\mathcal{N}}\alpha_{t}^{2}\sum_{k}\lambda_{t,k}^{a}(J_{C_{k}}(\pi_{f})-J_{C_{k}}(\theta_{t}))$
$\quad-\sum_{t\in\mathcal{N}}\alpha_{t}^{2}\sum_{i}\nu_{t,i}^{b}(J_{R_{i}}(\pi_{f})-J_{R_{i}}(\theta_{t}))$
$\leq D_{\mathrm{KL}}(\pi_{f}||\pi_{0})+\sum_{t}\alpha_{t}^{2}\underbrace{\left(\kappa_{1}+\frac{2R_{\mathrm{max}}\lambda_{\mathrm{max}}}{1-\gamma}\max(N,M)\right)}_{=:\kappa_{2}}.$
Since $\sum_{t\notin\mathcal{N}}\alpha_{t}\eta=\sum_{t}\alpha_{t}\eta-\sum_{t\in\mathcal{N}}\alpha_{t}\eta$,
$\sum_{t\in\mathcal{N}}\left(\alpha_{t}\left(\sum_{i}\nu_{t,i}^{a}(J_{R_{i}}(\pi_{f})-J_{R_{i}}(\theta_{t}))-\eta\right)\right)+\sum_{t}\alpha_{t}\eta\leq D_{\mathrm{KL}}(\pi_{f}||\pi_{0})+\sum_{t}\alpha_{t}^{2}\kappa_{2}.$
Due to the Robbins-Monro condition, the right term converges to real number. Since $\sum_{t}\alpha_{t}\eta=\infty$ in the left term, the following must be satisfied:
$\sum_{t\in\mathcal{N}}\left(\alpha_{t}\left(\sum_{i}\nu_{t,i}^{a}(J_{R_{i}}(\pi_{f})-J_{R_{i}}(\theta_{t}))-\eta\right)\right)=-\infty,$
which results in $\sum_{t\in\mathcal{N}}\alpha_{t}=\infty$. Now, let us define a preference $\omega$, where $\omega_{i}:=\bar{\nu}^{a}_{i}$. By replacing $\mu$ in (21) with $\pi_{\omega}^{*}$,
$\sum_{t\in\mathcal{N}}\left(\alpha_{t}\sum_{i}\nu_{t,i}^{a}(J_{R_{i}}(\pi_{\omega}^{*})-J_{R_{i}}(\theta_{t}))\right)-\sum_{t\notin\mathcal{N}}\left(\alpha_{t}\sum_{k}\lambda_{t,k}^{b}(J_{C_{k}}(\pi_{\omega}^{*})-J_{C_{k}}(\theta_{t}))\right)$
$\leq D_{\mathrm{KL}}(\pi_{\omega}^{*}||\pi_{0})+\kappa_{2}\sum_{t}\alpha_{t}^{2}.$
$\Rightarrow\sum_{t\in\mathcal{N}}\left(\alpha_{t}\sum_{i}\nu_{t,i}^{a}(J_{R_{i}}(\pi_{\omega}^{*})-J_{R_{i}}(\theta_{t}))\right)\leq D_{\mathrm{KL}}(\pi_{\omega}^{*}||\pi_{0})+\kappa_{2}\sum_{t}\alpha_{t}^{2}.$
Since the right term converges to a real number and $\sum_{t\in\mathcal{N}}\alpha_{t}=\infty$, the following must be satisfied:
$\lim_{t\to\infty}\sum_{i}\nu_{t,i}^{a}(J_{R_{i}}(\pi_{\omega}^{*})-J_{R_{i}}(\theta_{\mathcal{N}_{t}}))=0,$
where $\mathcal{N}_{t}$ is the $t$th element of $\mathcal{N}$. Consequently,
$\sum_{i}\bar{\nu}_{i}^{a}J_{R_{i}}(\pi_{\omega}^{*})=\sum_{i}\bar{\nu}_{i}^{a}\lim_{t\to\infty}J_{R_{i}}(\pi_{\mathcal{N}_{t}}),$
which means that $\lim_{t\to\infty}\pi_{\mathcal{N}_{t}}$ is also a LS optimal policy for the preference $\omega$. Due to Lemma B.1, the policy converges to a CP optimal policy. ∎
B.3 Proof of Theorem 4.3
See 4.3